four equal point charge of +3.0Cμ are placed in air at the four corners of a square that is 40cm apart on a side. Find the force on any of the charge


1
Expert's answer
2021-04-06T13:54:01-0400

Let's denote the corners of the square as ABCD (starting from the upper right corner). Let's the side of the square equals 0.4 meters. Let's find the electric force on charge "q_A" (which is located at corner A) due to other charges. There are three electric forces that acts on charge "q_A": the force "F_{BA}" directed upward, the force "F_{DA}" directed rightward and the force "F_{CA}" directed along the square diagonal at an angle of "45^{\circ}". Let's find the "x"- and "y"-components of this forces:


"F_{BA,x}=0, F_{BA,y}=\dfrac{kq_Bq_A}{r_{BA}^2},""F_{BA,y}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot(3\cdot10^{-6}\ C)^2}{(0.4\ m)^2}=0.51\ N,""F_{DA,x}=\dfrac{kq_Dq_A}{r_{DA}^2}, F_{DA,y}=0,""F_{DA,x}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot(3\cdot10^{-6}\ C)^2}{(0.4\ m)^2}=0.51\ N,""F_{CA,x}=\dfrac{kq_Cq_A}{r_{CA}^2}cos\theta,""F_{CA,x}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot(3\cdot10^{-6}\ C)^2}{(\sqrt{2}\cdot0.4\ m)^2}\cdot cos45^{\circ}=0.18\ N,""F_{CA,y}=\dfrac{kq_Cq_A}{r_{CA}^2}sin\theta,""F_{CA,y}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot(3\cdot10^{-6}\ C)^2}{(\sqrt{2}\cdot0.4\ m)^2}\cdot sin45^{\circ}=0.18\ N."

Then, we can write "x"- and "y"-components of resulting electric force that act on charge "q_A":


"F_x=F_{DA,x}+F_{CA,x}=0.51\ N+0.18\ N=0.69\ N,""F_y=F_{BA,y}+F_{CA,y}=0.51\ N+0.18\ N=0.69\ N."

Finally, we can find the resulting electric force that act on charge "q_A" from the Pythagorean theorem:


"F=\sqrt{F_x^2+F_y^2}=\sqrt{(0.69\ N)^2+(0.69\ N)^2}=0.97\ N."